A piece of metal with a mass of 97.6 grams was heated in a steam jacket to 94°C and then




plunged into 243.6 grams of water at 21 degrees * C causing the temperature to rise to 24 degrees * C What




is the specific heat of metal?

1
Expert's answer
2022-02-15T17:23:02-0500

"H =mc\Delta\theta"

Heat lost by metal

"= 97.6\u00d7c \u00d7 (94-24)= 6832c"

Heat gained by water

"243.6 \u00d74.184\u00d7(24-21)= 3057.7"


"6832c=3057.7\\x= 0.4476Jg^{-1}K^{-1}"